an investor buy 500 shares at $35 with 50% margin. After one year, the shares are sold for $50. The investor is subject to a 15% tax on gains, and inflation over the year was 4%. Ignoring interest and commissions, the investor's real after tax leveraged return is ?
Answer #1

Shares bought on margin means - it is the facility offered by the stock broker if you have $ 50 and bought stock with 50% margin means stock broker given you credit of $50 so you can purchase stock worth $100

Shares bought worth - 500 * $35 = $17500 means real investment = $17500/2 = $8750

Sold for $50

Profit ($50-$35)*500 = $7500 after tax return is $6375

therefore total return is $6375/$8750 * 100 = 72.86%

Real rate = (1+nominal rate) / (1+Inflation rate)

= 1.7286/1.04 = 1.4972

therefore real after tax return is 66.21%

Assuming Tax rate is applicable on sale of shares